# Task-driven Image Fusion with Learnable Fusion Loss

### v1

@@ Abstract
Multi-modal image fusion aggregates information from multiple sensor sources, achieving superior visual quality and perceptual characteristics compared to any single source, often enhancing downstream tasks. However, current fusion methods for downstream tasks still use predefined fusion objectives that potentially mismatch the downstream tasks, limiting adaptive guidance and reducing model flexibility. To address this, we propose Task-driven Image Fusion (TDFusion), a fusion framework incorporating a learnable fusion loss guided by task loss. Specifically, our fusion loss includes learnable parameters modeled by a neural network called the loss generation module. This module is supervised by the loss of downstream tasks in a meta-learning manner. The learning objective is to minimize the task loss of the fused images, once the fusion module has been optimized by the fusion loss. Iterative updates between the fusion module and the loss module ensure that the fusion network evolves toward minimizing task loss, guiding the fusion process toward the task objectives. TDFusion’s training relies solely on the loss of downstream tasks, making it adaptable to any specific task. It can be applied to any architecture of fusion and task networks. Experiments demonstrate TDFusion’s performance in both fusion and task-related applications, including four public fusion datasets, semantic segmentation, and object detection. The code will be released.
@@ 1 Introduction
Multi-modal image fusion [45, 37, 24, 63, 20, 66, 56] combines information from multiple sensors to produce a more holistic and detailed representation. Infrared images capture thermal radiation regardless of lighting conditions, while visible images provide more detail and texture. The advantages of fused images, such as increased information richness and improved robustness [31, 30, 14, 55, 54], make image fusion beneficial for downstream tasks. It often outperforms single-modal inputs in tasks such as semantic segmentation [9, 42, 26], object detection [4, 25], and others [15]. Most methods treat the fusion task as a simple image reconstruction problem, relying on predefined unsupervised loss functions [67, 60, 72, 73, 71] or perceptual loss [18, 16, 61]. These methods focus on predefined objectives for information aggregation, primarily achieving fusion at the visual level. Therefore, they often fail to capture crucial semantic information during feature extraction, hindering scene understanding and impairing task performance [10, 38, 19]. Recently, some fusion methods have further explored the mutual enhancement between fusion and downstream tasks. By cascading the fusion network with downstream task network [35, 28], the task loss constrains the fusion learning, ensuring the fused images satisfy the task requirements [42, 25]. Alternatively, some approaches embed high-level visual task features [62, 26, 60] or learn optimal initializations [29].
@@
However, although downstream tasks are integrated, existing frameworks still rely on predefined fusion loss terms that do not adapt dynamically to these tasks. The impact of downstream tasks, through specific combinations, remains limited. Manually defined losses inherently involve predefined guidance, often overlooking task-specific requirements. This guidance imposes manually designed prior constraints on the fusion process, limiting the dynamic and adaptive influence of downstream tasks on specific image pairs. These methods, whether embedding useful downstream task features [62, 26] or directly using downstream losses to guide fusion networks [42, 25], still face the limitations of fixed fusion loss terms. Networks tailored specifically for downstream tasks [60] introduce dependencies between fusion and tasks, restricting flexibility and limiting their applicability to various high-level vision tasks. To address these issues, we propose a downstream task-driven learning framework that includes a learnable loss. The fusion loss contains learnable parameters, generated by a loss generation module, and is designed to retain the intensity information of the source images for specific downstream tasks. The purpose of updating the fusion loss is to guide the fusion network in generating fused images that minimize downstream task loss, thereby enhancing adaptability. Moreover, the fusion loss update relies on the downstream task loss, making it independent of any specific task or network architecture.
@@
The loss generation module produces fusion losses, which are used to update fusion modules subsequently. This complexity presents challenges to standard end-to-end training. Fortunately, meta-learning techniques, which are strategies for learning how to learn, can effectively achieve the learning objectives of the loss generation module. This involves minimizing task losses for fused images through an optimized fusion loss. Meta-learning tackles common deep-learning challenges like limited data, high computational costs, and the need for better generalization. Key optimization areas include parameter tuning [7], optimization strategies [21], and designing network architectures [23]. In this paper, we draw inspiration from the Model-Agnostic Meta-Learning (MAML) approach [7] to train our loss generation module. Specifically, training the loss generation module involves two stages: inner update and outer update. During inner updates, the output of the loss generation module updates a surrogate fusion module without altering the original parameters. During outer updates, the fused image from the surrogate module is fed into the task network. The resulting task loss then updates the loss generation module through backpropagation. This alternating training ensures that the loss generation module consistently produces fusion losses that minimize the downstream task loss of the fused images.
@@
This paper introduces TDFusion, a task-oriented fusion framework driven by downstream tasks. It consists of a fusion module, a task module, and a loss generation module that learns to optimize the fusion loss. The fusion loss incorporates intensity preferences from source images and gradient preservation, guided by downstream task losses to refine intensity preferences. This model follows the general form of fusion loss used in advanced methods [67, 60, 72, 73], ensuring adaptability to various tasks. The updates of the loss generation module are performed using a meta-learning approach, optimizing the loss generation module parameters based on task losses from the fused images after each update of the fusion module. This process ensures that the loss function guides the fusion module continuously, optimizing feature aggregation and minimizing downstream task losses. The loss generation module dynamically adjusts through alternating updates with the fusion and task modules, generating optimal fusion losses at each model state. Our contributions can be summarized as follows:
@@
  • We propose TDFusion, a meta-learning-based fusion framework that leverages the loss function of downstream tasks for training. This method promotes task-driven fusion and alleviates challenges caused by the absence of ground truth in fused images. Moreover, this framework is agnostic of specific downstream tasks or network architectures, which enhances its adaptability and flexibility.
@@
  • We analyze the information preferences of downstream tasks such as semantic segmentation and object detection, providing deeper insights into multi-modal high-level tasks.
  • TDFusion achieves outstanding performance across fusion and high-level vision tasks, validated on four fusion datasets with semantic segmentation and object detection.
@@ 2 Related Work
2.1 DL-based Image Fusion
Deep learning-based image fusion methods have revolutionized the field by exploiting the powerful feature extraction capabilities of neural networks [49, 61]. These methods are broadly categorized into discriminative and generative approaches. Discriminative methods [65, 6, 69, 72] leverage the strong reconstruction ability of neural networks to directly learn the mapping between source and fused images [16, 18, 64, 66]. Generative methods, on the other hand, model the image generation process using generative approaches, integrating source images from a distributional perspective. These include methods based on Generative Adversarial Networks [32, 33, 34, 71] and diffusion models [68, 57]. Unified fusion methods [59, 73] bridge the gap between different fusion sub-tasks, incorporating strategies such as continual learning [51] and self-supervised decomposition techniques [22]. The introduction of registration modules helps to mitigate misalignment issues [12, 47, 52] in source images. Recent studies have further explored the synergy between fusion tasks and high-level vision tasks. These include leveraging downstream task losses to optimize fusion networks [42, 25], embedding high-level task features [26, 60, 62], and employing initialization techniques [29].
@@ 2.2 Meta-Learning in Vision
Meta-learning develops algorithms to automatically fine-tune hyperparameters for specific tasks, showing its versatility and effectiveness across various domains. MAML [7] and its variations [8, 36, 39] focus on learning efficient initialization parameters to quickly adapt to new tasks using minimal data. Meta-SGD [21] extends MAML by learning optimal update directions and rates, particularly beneficial in few-shot learning scenarios. Other methods like MW-Net [41] and L2RW [40] prioritize choosing relevant sample weights to tackle noisy data by using a compact validation set. Additionally, some studies focus on improving model adaptability through learning loss functions [11, 1, 3].
In image fusion, [17] learns filters to generate fused images at any resolution. MetaFusion [62] introduces a mechanism that improves image fusion and object detection by aligning semantics with fusion-specific features. Meta-learning also supports neural architecture search [27, 29] to identify optimal network architectures for image fusion and customizes network initialization for various tasks [29]. Furthermore, the continuous learning [50] integrates multi-task fusion is also in line with meta-learning principles.
@@ 3.1 Overview
Our TDFusion framework, as shown in Fig. 1,consists of a fusion network $\mathcal{F}(\cdot)$, a downstream task network $\mathcal{T}(\cdot)$, and a fusion loss generation module $\mathcal{G}(\cdot)$, which produces parameters for a learnable loss function. The parameters of these modules are denoted as $\theta_{\mathcal{F}}$, $\theta_{\mathcal{T}}$, and $\theta_{\mathcal{G}}$. The one-step updated clones of $\mathcal{F}$ and $\mathcal{T}$ are denoted as $\mathcal{F^{\prime}}$ and $\mathcal{T^{\prime}}$, with parameters $\theta_{\mathcal{F^{\prime}}}$ and $\theta_{\mathcal{T^{\prime}}}$. During the updates, the fusion network and the loss generation module alternate in learning, as depicted by blue and purple in Fig. 1. The update of the loss generation module consists of inner and outer updates, detailed in the following subsections. $\mathcal{L}_{f}$ and $\mathcal{L}_{t}$ represent the learnable fusion loss and task-specific loss, respectively, with their formulations provided in the next section.
@@ 3.4 Learning of loss generation module
Fig. 1 illustrates that the loss generation and fusion modules are trained alternately, ensuring the fusion loss is optimized at various stages of training and under different states of the fusion network. The process of training to optimize the fusion loss involves two key steps: the inner update and the outer update. In the inner update, clones of both the fusion network and the task network are generated, with each network undergoing a single training iteration using the fusion loss and task loss, respectively. This process is designed to obtain the state of the network guided by the fusion loss. During the outer update, the task loss of the fusion image produced by the updated clone is calculated. This loss is then backpropagated to the loss generation network. The goal of this step is to direct the fusion network to generate fusion images that result in lower downstream task losses, once guided by the fusion loss. The alternating updates between the inner and outer steps constitute the learning procedure for the loss generation module.
@@ 3.4.1 Inner Update
During the internal update phase, the fusion network $\mathcal{F}$ undergoes a single update guided by the fusion loss, which depends on the current state of network $\mathcal{G}$. This update primarily aims to compute the intermediate parameters $\theta_{\mathcal{F^{\prime}}}$ and $\theta_{\mathcal{R^{\prime}}}$, which are crucial for updating $\theta_{\mathcal{G}}$ in the subsequent phase. The upper section of the purple region in Fig. 1 illustrates this process. During this phase, the images from the meta-training set $\{I_{a}^{mtr},I_{b}^{mtr}\}$ are fed into the model, where $\mathcal{F}$ undergoes a single gradient descent update:
@@
$\theta_{\mathcal{G}}$ represent the parameters of the loss generation module $\mathcal{G}$, which determine the parameters for the learnable fusion loss. In the update equation, the step size $\eta_{\mathcal{F^{\prime}}}$ updates the fusion module. The module $\mathcal{F^{\prime}}$ temporarily substitutes $\mathcal{F}$, adjusting its parameters in one update step. Meanwhile, the parameters of $\mathcal{F}$, denoted as $\theta_{\mathcal{F}}$, remain unchanged. Similarly, $\mathcal{T^{\prime}}$ is updated in a single step using the parameters $\theta_{\mathcal{T}}$ from the current task network $\mathcal{T}$:
@@
The parameters $\theta_{\mathcal{F^{\prime}}}$ and $\theta_{\mathcal{T^{\prime}}}$ are updated during the inner update, which also ensures that the computation graph of $\theta_{\mathcal{F^{\prime}}}$ with respect to $\theta_{\mathcal{G}}$ is preserved. This preserved graph is essential for optimizing $\theta_{\mathcal{G}}$ during the outer update.
@@ 3.6 Network Architecture
TDFusion is composed of three modules: the fusion network, the downstream task network, and the loss generation network. The fusion network shares the same architecture as [2], a lightweight model built upon the Restormer Block (RTB) [58]. An adaptive fusion module is incorporated into this network to facilitate feature integration. Fig. 1 illustrates the structure of the loss generation module. It also employs the Restormer Block (RTB) [58] as core unit, receiving inputs $\{I_{a},I_{b}\}$. After applying $Softmax(\cdot)$, the final output guarantees $w_{a}^{ij}+w_{b}^{ij}=1$ for each pixel. This design ensures that the fused image meets the similarity constraints and removes the reliance on initialization within the loss generation module. The architecture of the downstream task network $\mathcal{T}(\cdot)$ depends on the specific. For learning of loss generation module, we chose the most lightweight models of SegFormer [5] and YOLOv8 [46] for semantic segmentation and object detection, respectively.

### v2

@@ Abstract
Multi-modal image fusion aggregates information from multiple sensor sources, achieving superior visual quality and perceptual features compared to single-source images, often improving downstream tasks. However, current fusion methods for downstream tasks still use predefined fusion objectives that potentially mismatch the downstream tasks, limiting adaptive guidance and reducing model flexibility. To address this, we propose Task-driven Image Fusion (TDFusion), a fusion framework incorporating a learnable fusion loss guided by task loss. Specifically, our fusion loss includes learnable parameters modeled by a neural network called the loss generation module. This module is supervised by the downstream task loss in a meta-learning manner. The learning objective is to minimize the task loss of fused images after optimizing the fusion module with the fusion loss. Iterative updates between the fusion module and the loss module ensure that the fusion network evolves toward minimizing task loss, guiding the fusion process toward the task objectives. TDFusion’s training relies entirely on the downstream task loss, making it adaptable to any specific task. It can be applied to any architecture of fusion and task networks. Experiments demonstrate TDFusion’s performance through fusion experiments conducted on four different datasets, in addition to evaluations on semantic segmentation and object detection tasks. The code is available at https://github.com/HaowenBai/TDFusion.
@@ 1 Introduction
Multi-modal image fusion [59, 51, 31, 79, 25, 82, 72] combines information from multiple sensors to produce a more holistic and detailed representation. Infrared images capture thermal radiation regardless of lighting conditions, while visible images provide richer texture details. Fused images enhance downstream tasks through improved information density and robustness [45, 41, 18, 71, 70, 62, 63], outperforming single-modal inputs in semantic segmentation [13, 56, 34, 28], object detection [8, 32], and other related applications [3, 20, 27, 1]. Conventional methods typically treat fusion as image restoration using unsupervised loss [83, 76, 88, 89, 87] or perceptual loss [23, 21, 77]. These approaches prioritize visual-level fusion through predefined aggregation objectives, often neglecting semantic feature extraction. This limitation hinders scene interpretation and task performance [14, 52, 24, 64]. Recent advances explore the mutual enhancement between fusion and downstream tasks [37]. By cascading the fusion network with downstream task network [49, 39], the task loss constrains the fusion learning, ensuring the fused images meet the task requirements [56, 32]. Alternatively, some methods incorporate high-level visual task features [78, 34, 76] or focus on learning optimal initializations [40] to enhance fusion.
@@
While integrating downstream tasks, existing frameworks still rely on predefined fusion loss terms lacking dynamic adaptation. The impact of downstream tasks remains limited due to specific combinations. Manually defined losses preserve predefined guidance, frequently overlooking task-specific requirements. This guidance imposes manually designed prior constraints on the fusion process, limiting the dynamic and adaptive influence of downstream tasks on specific image pairs. These approaches, whether incorporating task features [78, 34] or employing task losses [56, 32], still face the limitations of fixed fusion loss terms. Task-specialized networks [76] create fusion-task dependencies, restricting flexibility and limiting their applicability to various high-level vision tasks. We address these limitations through a task-driven framework with learnable loss. The fusion loss contains learnable parameters, generated by a loss generation module, and is designed to retain the intensity information of the source images for specific downstream tasks. The purpose of updating the fusion loss is to guide the fusion network in generating fused images that minimize downstream task loss, thereby enhancing adaptability. Moreover, the fusion loss update relies on the downstream task loss, making it independent of any specific task or network architecture.
@@
The loss generation module produces fusion losses for subsequent fusion module updates. This complexity presents challenges to standard end-to-end training. Fortunately, meta-learning techniques, which are strategies for learning how to learn, can effectively achieve the learning objectives of the loss generation module. This involves minimizing task loss for fused images through an optimized fusion loss. Meta-learning tackles common deep-learning challenges like limited data, high computational costs, and the need for better generalization. Core optimization areas include parameter tuning [11], optimization strategies [26], and network architectures search [30]. In this paper, we draw inspiration from the Model-Agnostic Meta-Learning (MAML) approach [11] to train our loss generation module. Specifically, training the loss generation module involves two stages: inner updates and outer updates. During inner updates, the output of the loss generation module updates a surrogate fusion module without altering the original parameters. During outer updates, the fused image from the surrogate module is fed into the task network. The resulting task loss then updates the loss generation module through backpropagation. This alternating training ensures that the loss generation module consistently produces fusion losses that minimize the downstream task loss of the fused images.
@@
This paper introduces TDFusion, a task-oriented fusion framework driven by downstream tasks. It consists of a fusion module, a task module, and a loss generation module that learns to optimize the fusion loss. The fusion loss incorporates intensity preferences from source images and gradient preservation, guided by downstream task loss to refine intensity preferences. This model follows the general form of fusion loss used in advanced methods [83, 76, 88, 89], ensuring adaptability to various tasks. The updates of the loss generation module are performed using a meta-learning approach, optimizing the loss generation module parameters based on task loss from the fused images after each update of the fusion module. This process ensures that the loss function guides the fusion module continuously, optimizing feature aggregation and minimizing downstream task losses. The loss generation module dynamically adjusts through alternating updates with the fusion and task modules, generating optimal fusion losses at each model state.
Our contributions can be summarized as follows:
@@
  • We propose TDFusion, a meta-learning-based fusion framework that leverages the loss functions of downstream tasks for training. This method promotes task-driven fusion and alleviates challenges caused by the absence of ground truth. Moreover, this framework is agnostic of specific downstream tasks or network architectures, which enhances its adaptability and flexibility.
@@
  • We analyze the information preferences of downstream tasks such as semantic segmentation and object detection, providing deeper insights into multi-modal high-level vision tasks.
  • TDFusion achieves outstanding performance in both fusion and high-level vision tasks, validated on four fusion datasets with semantic segmentation and object detection.
@@ 2 Related Work
2.1 Deep learning-based Image Fusion
Deep learning-based image fusion methods have revolutionized the field by exploiting the powerful feature extraction capabilities of neural networks [42, 43, 44, 33, 36, 35, 6, 65, 77, 4]. These methods are broadly categorized into discriminative and generative approaches. Discriminative methods [81, 10, 85, 88] leverage the strong reconstruction ability of neural networks to directly learn the mapping between source and fused images [21, 23, 80, 82]. Generative methods, on the other hand, model the image generation process using generative approaches, integrating source images from a distributional perspective. These include methods based on Generative Adversarial Networks [46, 47, 48, 87] and diffusion models [84, 73]. Unified fusion methods [66, 75, 89] bridge the gap between different fusion sub-tasks, incorporating strategies such as continual learning [67] and self-supervised decomposition techniques [29]. The introduction of registration modules helps to mitigate misalignment issues [16, 60, 68] in source images. Recent studies have further explored the synergy between fusion and high-level vision tasks. These include leveraging downstream task losses to optimize fusion networks [56, 32], embedding high-level task features [34, 76, 78], and employing initialization techniques [40].
@@ 2.2 Meta-Learning in Vision
Meta-learning develops algorithms to automatically fine-tune hyperparameters for specific tasks, showing its versatility and effectiveness across various domains. MAML [11] and its variations [12, 50, 53] focus on learning efficient initialization parameters to quickly adapt to new tasks using minimal data. Meta-SGD [26] extends MAML by learning optimal update directions and rates, beneficial in few-shot learning scenarios. Other approaches like MW-Net [55] and L2RW [54] emphasize selecting relevant sample weights to tackle noisy data using a compact validation set. Additionally, some studies focus on improving model adaptability through learning loss functions [15, 2, 7].
In image fusion, learnable filters [22] enable the fusion of images at arbitrary resolutions. MetaFusion [78] introduces a mechanism that improves image fusion and object detection by aligning semantics with fusion-specific features. ReFusion [5] guides the learnable fusion loss for various fusion tasks by reconstructing the source images. Meta-learning also supports neural architecture search [38, 40] to identify optimal network architectures for image fusion and customizes network initialization for various tasks [40].
@@ 3.1 Overview
Our TDFusion framework, as shown in Fig. 1, consists of a fusion network $\mathcal{F}(\cdot)$, a downstream task network $\mathcal{T}(\cdot)$, and a fusion loss generation module $\mathcal{G}(\cdot)$, which produces parameters for a learnable loss function. The parameters of these modules are denoted as $\theta_{\mathcal{F}}$, $\theta_{\mathcal{T}}$, and $\theta_{\mathcal{G}}$, respectively. The one-step updated clones of $\mathcal{F}$ and $\mathcal{T}$ are denoted as $\mathcal{F^{\prime}}$ and $\mathcal{T^{\prime}}$, with parameters $\theta_{\mathcal{F^{\prime}}}$ and $\theta_{\mathcal{T^{\prime}}}$. During the updates, the fusion network and the loss generation module alternate in learning, as depicted by blue and purple in Fig. 1. The update of the loss generation module consists of inner and outer updates, detailed in the following subsections. $\mathcal{L}_{f}$ and $\mathcal{L}_{t}$ represent the learnable fusion loss and task-specific loss, with their formulations provided in the next section.
@@ 3.4 Learning of loss generation module
Fig. 1 illustrates that the loss generation and fusion modules are trained alternately, ensuring the fusion loss is optimized at various stages of training and under different states of the fusion network. The process of training to optimize the fusion loss involves two key steps: the inner update and the outer update. In the inner update, clones of both the fusion network and the task network are generated, with each network undergoing a single training iteration using the fusion loss and task loss, respectively. This process is designed to obtain the state of the network guided by the fusion loss. During the outer update, the task loss of the fusion image produced by the updated clone is calculated. This loss is then backpropagated to the loss generation network. The goal of this step is to direct the fusion network to generate fusion images that result in lower downstream task loss, once guided by the fusion loss. The alternating updates between the inner and outer steps constitute the learning procedure for the loss generation module.
@@ 3.4.1 Inner Update
During the inner update phase, the fusion network $\mathcal{F}$ undergoes a single update guided by the fusion loss, which depends on the current state of network $\mathcal{G}$. This update primarily aims to compute the intermediate parameters $\theta_{\mathcal{F^{\prime}}}$ and $\theta_{\mathcal{R^{\prime}}}$, which are crucial for updating $\theta_{\mathcal{G}}$ in the subsequent phase. The upper section of the purple region in Fig. 1 illustrates this process. During this phase, the images from the meta-training set $\{I_{a}^{mtr},I_{b}^{mtr}\}$ are fed into the model:
@@
where $\mathcal{F}$ undergoes a single gradient descent update. $\theta_{\mathcal{G}}$ represent the parameters of the loss generation module $\mathcal{G}$, which determine the parameters for the learnable fusion loss. And the notation $\eta_{\mathcal{F^{\prime}}}$ refers to the step size. The module $\mathcal{F^{\prime}}$ temporarily substitutes $\mathcal{F}$, adjusting its parameters in one update step. Meanwhile, the parameters of $\mathcal{F}$, denoted as $\theta_{\mathcal{F}}$, remain unchanged. Similarly, $\mathcal{T^{\prime}}$ is updated in a single step using the parameters $\theta_{\mathcal{T}}$ from the current task network $\mathcal{T}$:
@@
The parameters $\theta_{\mathcal{F^{\prime}}}$ and $\theta_{\mathcal{T^{\prime}}}$ are both updated during the inner update, which also ensures that the computation graph of $\theta_{\mathcal{F^{\prime}}}$ with respect to $\theta_{\mathcal{G}}$ is preserved. This preserved graph is essential for optimizing $\theta_{\mathcal{G}}$ during the outer update.
@@ 3.5 Learning of Fusion Network
Algorithm 1 TDFusion Training Algorithm
1:Require: Training set $\{I_{a}^{ftr},I_{b}^{ftr}\}$ with size N.
2:Output: Thoroughly trained $\theta_{\mathcal{F}}$, $\theta_{\mathcal{T}}$, $\theta_{\mathcal{G}}$.
4:for $epoch=1$ to L do
5:     Sample $\{I_{a}^{mtr},I_{b}^{mtr}\}$ and $\{I_{a}^{mts},I_{b}^{mts}\}$.
6:     for $step=1$ to M do
7:   % Inner update: apply $\mathcal{G}$.
8:         Sample $(I_{a}^{mtr},I_{b}^{mtr})$ and get $(I_{f}^{mtr},y^{mtr})$.
10:   % Outer update: optimize $\mathcal{G}$.
11:         Sample $(I_{a}^{mts},I_{b}^{mts})$ and get $(I_{f}^{mts},y^{mts})$.
12:         Update $\theta_{\mathcal{G}}$ by Eq. 6.
13:     end for
14:     for $step=1$ to N do
15:   % Fusion update: optimize $\mathcal{F}$ and $\mathcal{T}$.
16:         Sample $(I_{a}^{ftr},I_{b}^{ftr})$ and get $(I_{f}^{ftr},y^{ftr})$.
17:         Update $\theta_{\mathcal{F}}$ and $\theta_{\mathcal{T}}$ by Eq. 8 and Eq. 9.
18:     end for
19:end for
@@ 3.6 Network Architecture
TDFusion is composed of three modules: the fusion network, the downstream task network, and the loss generation network. The fusion network shares the same architecture as [5], a lightweight model built upon the Restormer Block (RTB) [74]. An adaptive fusion module is incorporated into this network to facilitate feature integration. Fig. 1 illustrates the structure of the loss generation module. It also employs the Restormer Block (RTB) [74] as its primary component, receiving inputs $\{I_{a},I_{b}\}$. After applying $Softmax(\cdot)$, the final output guarantees $w_{a}^{ij}+w_{b}^{ij}=1$ for each pixel. This design ensures that the fused image meets the similarity constraints and removes the reliance on initialization within the loss generation module. The architecture of the downstream task network $\mathcal{T}(\cdot)$ depends on the specific task. For learning of loss generation module, we chose the most lightweight models of SegFormer [9] and YOLOv8 [19] for semantic segmentation and object detection, respectively.